A student opens his laptop with a 2.5 N force at an angle of 70 degrees against the horizontal. Find the torque on the cover of the laptop if the laptop is 20 cm wide.*


1
Expert's answer
2022-04-02T10:10:37-0400

"M=Fr\sin\alpha=0.47~N\cdot m ."